# Stable Trajectory Clustering: An Efficient Split and Merge Algorithm

### v1

@@ VI Empirical Results
VI-A DataSets and Environment Setup
@@
We use two kinds of real trajectory datasets. Dataset 1 is related to a four-way traffic intersection, including maneuvers such as turns and U-turns, and contains 1900 trajectories [48]. Dataset 2, which is known as the Vehicle Energy Dataset (VED), includes information from 383 privately-owned cars in Ann Arbor, Michigan, USA  [49]. This Dataset contains the trajectories of cars driving on both highways and traffic-dense areas from November 2017 to November 2018 and was collected by onboard OBD-II loggers. All algorithms were implemented in Matlab 2020 and executed on a laptop with an Intel Core(TM) i7-4510U processor and 12GB of memory. Before running the algorithms, the data were preprocessed by removing duplicated spatial coordinates and standardizing all time intervals.
VI-B Effectiveness and Parameter Sensitivity
@@
First, we will consider ranges of values for the two user-specified DBSCAN parameters, $\varepsilon$, and $MinLns$. For each set of DBSCAN parameters, we then apply the whole-trajectory clustering algorithms on Dataset 1. In sub-trajectory clustering, we further explore the effects of the sliding window parameters $S$ and $W$ on Dataset 1. Following that, we execute the stable trajectory clustering algorithm and analyze its effectiveness in Dataset 2. Sensitivity analysis for DBSCAN line segment clustering involves evaluating the number of clusters and cluster density across the ranges of user-specified input parameters. These two parameters can be tuned depending on the characteristics of the data, and as mentioned before, their values impact the clustering result. Specifically, larger values of $\varepsilon$ result in larger neighborhoods, which tend to produce larger and less compact clusters. In contrast, the value of $MinLns$ plays a significant role in assigning core or border status to line segments. Higher values require more neighboring line segments to classify a segment as core, potentially reducing the number of clusters or affecting cluster boundaries. Fig. 5 aims to provide a general overview of the impact of these two parameters on whole-trajectory clustering for Dataset 1. As expected, the number of clusters changes as a function of the values of $\varepsilon$ and $MinLns$. A similar analysis can be conducted to examine the influence of other values for these parameters, as illustrated in Fig. 6.
Figure 5: Effect of $\varepsilon$ and $MinLns$ on the number of clusters.
@@
Figure 6: Whole-trajectory clustering on Dataset 1. Increasing $\varepsilon$ (from 8 to 10) and $MinLns$ (from 5 to 80) leads to denser clusters, alters their count, and impacts the number of outlier trajectories. These changes stem from line segment clustering in each time interval, impacting the overall whole-trajectory clustering.
@@
Figure 7: Sub-trajectory clustering on Dataset 1 during the second iteration of the algorithm. (a) Comparison of sub-trajectory clustering results using different step sizes ($S=3$ and $S=11$) with a fixed window size ($W=15$). (b) Comparison of sub-trajectory clustering results using different window sizes ($W=15$ and $W=21$) with a fixed step size ($S=5$).
@@
Fig. 7 demonstrates the effects of varying $S$ and $W$ on sub-trajectory clustering for Dataset 1, which contains trajectories divided into 50-time intervals. The results correspond to the second iteration of the algorithm. For a fixed window size $W$, increasing the parameter $S$ causes the sliding window to move forward more significantly during the same iterations of the algorithm (7(a)). On the other hand, the larger the sliding window size, the more extended time intervals it encompasses, leading to a wider range of sub-trajectories (7(b)). The execution of the whole-trajectory clustering algorithm on Dataset 2 is illustrated in Fig. 8. For further clarity, magnified views within the figures are provided, highlighting parts of cluster $C_{6}$ and $C_{9}$, along with their outlier trajectories in detail. After applying the stable trajectory clustering algorithm, the minor deviations of the outlier trajectory are disregarded, and the trajectories are reassigned to their cluster (8(b)).
@@
Figure 8: The stable trajectory clustering algorithm results on Dataset 2 over 101-time intervals. (a) The whole-trajectory clustering result before running the stable trajectory clustering algorithm ($\varepsilon=0.006$ and $MinLns=2$). The magnified view of $C_{6}$ and $C_{9}$ includes outlier trajectories that temporarily split from their cluster for a few seconds. (b) The whole trajectory clustering result after running the stable trajectory clustering algorithm. The magnified view shows that former outlier trajectories are added to $C_{6}$ and $C_{9}$.
@@
As discussed in section V, in the stable trajectory clustering algorithm, the number of time intervals and the deviation amount in these intervals are important factors in deciding whether or not to ignore anomalies. Based on these two factors, $\mu_{min}$ is calculated for each of 12 clusters and their corresponding outlier trajectories. For example, this value for cluster $C_{6}$ and its 9 outlier trajectories is $0.0074$. Out of these, 4 are correctly assigned to their respective clusters after applying the stable clustering algorithm. Fig. 9 illustrates the temporal distance of outlier trajectories to their nearest cluster members, recognized as the distance from cluster $C_{6}$. Each box plot represents the distribution of the distances between the segments of an outlier trajectory and the segments of its closest trajectory within the cluster across 101-time intervals. As is clear, outlier trajectories with a mean less than $\mu$ are identified as non-outlier trajectories. This means that the stable trajectory clustering algorithm can ignore insignificant anomalies in a few time intervals. These time intervals are shown in Fig. 10. This figure depicts each outlier trajectory as a series of 101 small rectangles, each corresponding to a time interval. It illustrates not only the time intervals during which each outlier trajectory remained within cluster $C_{6}$ but also those during which it was temporarily split from the cluster.
@@
Figure 9: Temporal distance of outlier trajectories from cluster $C_{6}$. In these box plots, each line represents the average distances over 101-time intervals. The dashed horizontal line indicates $\mu_{min}$, a threshold value calculated to indicate the maximum acceptable average distance for identifying non-outlier trajectories.
@@
Figure 10: Temporal membership of outlier trajectories in cluster $C_{6}$ over 101-time intervals. Blue rectangles represent the time intervals during which the outlier trajectory remains within the cluster, while pink rectangles signify the time intervals when it is outside the cluster. Non-outlier trajectories are represented by the darker blue and darker pink areas, based on indicated non-outlier trajectories in Fig. 9. The darker pink regions are also hatched, indicating time intervals where outlier trajectory deviations were disregarded.
@@
As mentioned before, VED (Dataset 2) provides a rich and diverse collection of real-world driving data, including GPS trajectories and energy consumption records, offering an invaluable resource for analyzing driving behavior and energy use. Although the stable clustering algorithm on the VED data trajectories can detect minor deviations in outlier trajectories, it considers those with sharp changes in direction, which are highly noticeable, as real outlier trajectories. This process improves the quality and reliability of the clustering results, as it considers driving patterns that reflect typical behavior. For example, a trajectory that was mistakenly labeled as an outlier due to a brief stop or a slight turn can now be correctly included in its cluster, improving the representation of real driving behavior. The improved clusters lead to more accurate energy consumption analyses and make the analysis more reliable for real-world applications like eco-driving strategies, trajectory optimization, and traffic flow studies. This filtering slightly reduces the average Silhouette coefficient score. This metric quantifies how well data trajectories fit within their clusters (cohesion) and how distinctly they are separated from others (separation). This reduction is due to assigning some outlier trajectories to their respective clusters after applying the stable trajectory clustering algorithm. The Silhouette score is calculated for Dataset 1 and Dataset 2 using different DBSCAN parameter settings before and after executing the stable trajectory clustering algorithm, as shown in Table I. In this table, in addition to this metric, the standard deviation is also computed to indicate the dispersion of the Silhouette scores of trajectories and the range within which these values vary. The key aspect for comparison is the difference in the Silhouette score before and after executing the algorithm. The smaller this difference, and the more trajectories successfully assigned to their respective clusters, the more effective the algorithm is. A lower difference indicates that the algorithm has successfully identified outlier trajectories with minor deviations, thus preserving the overall clustering structure and maintaining the average Silhouette score with minimal change. It should be noted that an outlier trajectory may either connect at least two clusters or be assigned to a single cluster by the stable trajectory clustering algorithm. This occurs because different segments of the outlier trajectory may belong to different clusters. After running the algorithm, the Silhouette score is evaluated for both cases, and the outlier trajectory is ultimately placed where it results in the smallest decrease in the Silhouette score.
@@
#1 $2$ $0.0075$ $13$ $13$ $0.640\pm 0.19$ $13$ $6$ $0.607\pm 0.20$
$2$ $0.0060$ $12$ $18$ $0.653\pm 0.21$ $12$ $7$ $0.617\pm 0.22$
$3$ $0.0065$ $11$ $17$ $0.661\pm 0.19$ $11$ $9$ $0.623\pm 0.20$
#2 $8$ $5$ $18$ $121$ $0.709\pm 0.19$ $18$ $6$ $0.702\pm 0.19$
$10$ $20$ $15$ $42$ $0.750\pm 0.12$ $15$ 00 $0.747\pm 0.12$
$60$ $80$ $11$ $40$ $0.773\pm 0.08$ $11$ 00 $0.722\pm 0.17$
TABLE I: Comparison of clustering results before and after applying the stable trajectory clustering (STC) algorithm for two datasets. The results show how STC improves the cluster structure as it assigns outlier trajectories (Out.) to their respective clusters, which causes a slight reduction in the Silhouette score. A smaller difference in the Silhouette score before and after STC indicates higher effectiveness.
Table II demonstrates the runtime performance of trajectory clustering approaches, where simpler datasets and lower parameter settings yield faster results, while more complex datasets and higher parameter settings require significantly more computation time.
@@
TABLE II: Runtime performance comparison of trajectory clustering algorithms: Split-Merge (with $\varepsilon$, MinLns), Whole-Trajectory, and Sub-Trajectory methods with parameters $S$ and $W$. Times reported in seconds for Datasets 1 and 2.
VII Conclusion
@@
Trajectory clustering algorithms can analyze either whole trajectories or subsets of them. In this paper, we proposed both a whole-trajectory clustering algorithm and a sub-trajectory clustering algorithm. To discover the continuous object movements in each time interval, the split and merge procedures of line segments were implemented via DBSCAN line segment clustering. We employed the average Euclidean distance to compute the distance between the line segments. In sub-trajectory clustering, we employed the sliding window model to recognize similar ranges of sub-trajectories. We introduced a novel algorithm, called the stable trajectory clustering algorithm, that represents an advanced leap forward in the literature on trajectory clustering. This algorithm addresses the situation when a moving object splits from its cluster for some time intervals and again joins. Ignoring these minor anomalies on large scales assists analysts in performing trajectory clustering without being distracted by insignificant deviations, maintains the integrity of the data, and improves the precision of clustering results. For instance, vehicles stopping at traffic lights, pedestrians pausing at shop windows, or animals detouring briefly during migration should still be considered part of the main movement cluster. To determine whether the anomalies of outlier trajectories are real or temporary, the number of time intervals in which deviations occur and the magnitudes of those deviations are important factors. Thus, we used the mean absolute deviation concept to analyze these factors. The empirical results on real data illustrated the effectiveness of the algorithm in detecting anomalies. Furthermore, the Silhouette score was used to evaluate the quality of the clustering results before and after applying the algorithm. The comparison showed that the effectiveness and robustness of the algorithm increase when two conditions are met: (1) the difference in Silhouette scores before and after applying the algorithm is minimized, and (2) a greater number of outlier trajectories with minor anomalies are correctly reassigned to their respective clusters. The proposed algorithms exhibit sensitivity to the input parameters $\varepsilon$ and $MinLns$, with optimal values being highly dependent on the characteristics of the moving data. This represents a limitation of our approach and warrants further investigation in future research. Additionally, exploring alternative concepts or methods to assess the stable trajectory clustering, beyond the mean absolute deviation approach used in this study, could provide valuable insights.

### v2

@@ 6. Empirical Results
6.1. DataSets and Environment Setup
@@
We use two kinds of real trajectory datasets. Dataset 1 is related to a four-way traffic intersection, including maneuvers such as turns and U-turns, and contains 1900 trajectories, organized into 19 ground-truth clusters (57). We selected this dataset because the ground truth is necessary to evaluate the impact of the stable trajectory clustering algorithm against outlier trajectories. Dataset 2, which is known as the Vehicle Energy Dataset (VED), includes information from 383 privately-owned cars in Ann Arbor, Michigan, USA  (44). This Dataset contains the trajectories of cars driving on both highways and traffic-dense areas from November 2017 to November 2018 and was collected by onboard OBD-II loggers. Due to the difficulty of demonstrating the impact of the stable trajectory clustering algorithm on dense datasets, we selected this dataset to allow for a clear visualization of cluster quality.
All algorithms were implemented in MATLAB 2020 and executed on a laptop with an Intel Core(TM) i7-4510U processor and 12GB of memory. Before running the algorithms, the data were preprocessed by removing duplicated spatial coordinates and adjusting all trajectories to a uniform length by linear interpolation, which was fundamental for establishing a common temporal framework for our analysis.
6.2. Effectiveness and Parameter Sensitivity
First, we will consider ranges of values for the two user-specified DBSCAN parameters, $\varepsilon$, and $MinLns$. For each set of DBSCAN parameters, we then apply the whole-trajectory clustering algorithms on Dataset 1. In sub-trajectory clustering, we further explore the effects of the sliding window parameters $S$ and $W$ on Dataset 1. Following that, we execute the stable trajectory clustering algorithm and analyze its effectiveness in Dataset 2.
@@ 6.2. Effectiveness and Parameter Sensitivity
Sensitivity analysis for DBSCAN line segment clustering involves evaluating the number of clusters and cluster density across the ranges of user-specified input parameters. These two parameters can be tuned depending on the characteristics of the data, and as mentioned before, their values impact the clustering result. Specifically, larger values of $\varepsilon$ result in larger neighborhoods, which tend to produce larger and less compact clusters. In contrast, the value of $MinLns$ plays a significant role in assigning core or border status to line segments. Higher values require more neighboring line segments to classify a segment as core, potentially reducing the number of clusters or affecting cluster boundaries. Fig. 5 aims to provide a general overview of the impact of these two parameters on whole-trajectory clustering for Dataset 1. As expected, the number of clusters changes as a function of the values of $\varepsilon$ and $MinLns$. A similar analysis can be conducted to examine the influence of other values for these parameters, as illustrated in Fig. 6.
Figure 5. Effect of $\varepsilon$ and $MinLns$ on the number of clusters.
@@
Figure 6. Whole-trajectory clustering on Dataset 1. Increasing $\varepsilon$ (from 8 to 10) and MinLnsMinLns (from 5 to 80) leads to denser clusters, alters their count, and impacts the number of outlier trajectories. These changes stem from line segment clustering in each time interval, impacting the overall whole-trajectory clustering.
@@
Figure 7. Sub-trajectory clustering on Dataset 1 during the second iteration of the algorithm. (a) Comparison of sub-trajectory clustering results using different step sizes ($S=3$ and $S=11$) with a fixed window size ($W=15$). (b) Comparison of sub-trajectory clustering results using different window sizes ($W=15$ and $W=21$) with a fixed step size ($S=5$).
@@
Fig. 7 demonstrates the effects of varying $S$ and $W$ on sub-trajectory clustering for Dataset 1, which contains trajectories divided into 50-time intervals. The results correspond to the second iteration of the algorithm. For a fixed window size $W$, increasing the parameter $S$ causes the sliding window to move forward more significantly during the same iterations of the algorithm (7(a)). On the other hand, the larger the sliding window size, the more extended time intervals it encompasses, leading to a wider range of sub-trajectories (7(b)).
The execution of the whole-trajectory clustering algorithm on Dataset 2 is illustrated in Fig. 8. For further clarity, magnified views within the figures are provided, highlighting parts of cluster $C_{6}$ and $C_{9}$, along with their outlier trajectories in detail. After applying the stable trajectory clustering algorithm, the minor deviations of the outlier trajectory are disregarded, and the trajectories are reassigned to their cluster (8(b)).
@@
Figure 8. The stable trajectory clustering algorithm results on Dataset 2 over 101 time intervals. (a) The whole-trajectory clustering result before running the stable trajectory clustering algorithm ($\varepsilon=0.006$ and $MinLns=2$). The magnified view of $C_{6}$ and $C_{9}$ includes outlier trajectories that temporarily split from their cluster for a few seconds. (b) The whole trajectory clustering result after running the stable trajectory clustering algorithm. The magnified view shows that former outlier trajectories are added to $C_{6}$ and $C_{9}$.
@@
As discussed in section 5, in the stable trajectory clustering algorithm, the number of time intervals and the deviation amount in these intervals are important factors in deciding whether or not to ignore anomalies. Based on these two factors, $\mu_{min}$ is calculated for each of 12 clusters and their corresponding outlier trajectories. For example, this value for cluster $C_{6}$ and its 9 outlier trajectories is $0.0074$. Out of these, 4 are correctly assigned to their respective clusters after applying the stable clustering algorithm. Fig. 9 illustrates the temporal distance of outlier trajectories to their nearest cluster members, recognized as the distance from cluster $C_{6}$. Each box plot represents the distribution of the distances between the segments of an outlier trajectory and the segments of its closest trajectory within the cluster across 101 time intervals. As is clear, outlier trajectories with a mean less than $\mu$ are identified as non-outlier trajectories. This means that the stable trajectory clustering algorithm can ignore insignificant anomalies in a few time intervals. These time intervals are shown in Fig. 10. This figure depicts each outlier trajectory as a series of 101 small rectangles, each corresponding to a time interval. It illustrates not only the time intervals during which each outlier trajectory remained within cluster $C_{6}$ but also those during which it was temporarily split from the cluster.
@@
Figure 9. Temporal distance of outlier trajectories from cluster $C_{6}$. In these box plots, each line represents the average distances over 101 time intervals. The dashed horizontal line indicates $\mu_{min}$, a threshold value calculated to indicate the maximum acceptable average distance for identifying non-outlier trajectories.
@@
Figure 10. Temporal membership of outlier trajectories in cluster $C_{6}$ over 101-time intervals. Blue rectangles represent the time intervals during which the outlier trajectory remains within the cluster, while pink rectangles signify the time intervals when it is outside the cluster. Non-outlier trajectories are represented by the darker blue and darker pink areas, based on the indicated non-outlier trajectories in Fig. 9. The darker pink regions are also hatched, indicating time intervals where outlier trajectory deviations were disregarded.
@@
As mentioned before, VED (Dataset 2) provides a rich and diverse collection of real-world driving data, including GPS trajectories and energy consumption records, offering an invaluable resource for analyzing driving behavior and energy use. Although the stable clustering algorithm on the VED data trajectories can detect minor deviations in outlier trajectories, it considers those with sharp changes in direction, which are highly noticeable, as real outlier trajectories. This process improves the quality and reliability of the clustering results, as it considers driving patterns that reflect typical behavior. For example, a trajectory that was mistakenly labeled as an outlier due to a brief stop or a slight turn can now be correctly included in its cluster, improving the representation of real driving behavior. The improved clusters lead to more accurate energy consumption analyses and make the analysis more reliable for real-world applications like eco-driving strategies, trajectory optimization, and traffic flow studies. This filtering slightly reduces the average Silhouette coefficient score. This metric quantifies how well data trajectories fit within their clusters (cohesion) and how distinctly they are separated from others (separation). This reduction is due to assigning some outlier trajectories to their respective clusters after applying the stable trajectory clustering algorithm. The Silhouette score is calculated for Dataset 1 and Dataset 2 using different DBSCAN parameter settings before and after executing the stable trajectory clustering algorithm, as shown in Table 1. In this table, in addition to this metric, the standard deviation is also computed to indicate the dispersion of the Silhouette scores of trajectories and the range within which these values vary. The key aspect for comparison is the difference in the Silhouette score before and after executing the algorithm. The smaller this difference, and the more trajectories successfully assigned to their respective clusters, the more effective the algorithm is. A lower difference indicates that the algorithm has successfully identified outlier trajectories with minor deviations, thus preserving the overall clustering structure and maintaining the average Silhouette score with minimal change. It should be noted that an outlier trajectory may either connect at least two clusters or be assigned to a single cluster by the stable trajectory clustering algorithm. This occurs because different segments of the outlier trajectory may belong to different clusters. After running the algorithm, the Silhouette score is evaluated for both cases, and the outlier trajectory is ultimately placed where it results in the smallest decrease in the Silhouette score.
@@
#1 22 $0.0075$ $13$ $13$ $0.640\pm 0.19$ $13$ 66 $0.607\pm 0.20$
22 $0.0060$ $12$ $18$ $0.653\pm 0.21$ $12$ 77 $0.617\pm 0.22$
33 $0.0065$ $11$ $17$ $0.661\pm 0.19$ $11$ 99 $0.623\pm 0.20$
#2 88 55 $18$ $121$ $0.709\pm 0.19$ $18$ 66 $0.702\pm 0.19$
$10$ 2020 $15$ $42$ $0.750\pm 0.12$ $15$ 0 $0.747\pm 0.12$
$60$ $80$ $11$ $40$ $0.773\pm 0.08$ $11$ 0 $0.722\pm 0.17$
Table 1. Comparison of clustering results before and after applying the stable trajectory clustering (STC) algorithm for two datasets. The results show how STC improves the cluster structure as it assigns outlier trajectories (Out.) to their respective clusters, which causes a slight reduction in the Silhouette score. A smaller difference in the Silhouette score before and after STC indicates higher effectiveness.
Table 2 demonstrates the runtime performance of trajectory clustering approaches, where simpler datasets and lower parameter settings yield faster results, while more complex datasets and higher parameter settings require significantly more computation time.
@@
Table 2. Runtime performance comparison of trajectory clustering algorithms: Split-Merge (with $\varepsilon$, MinLns), Whole-Trajectory, and Sub-Trajectory methods with parameters $S$ and $W$. Times reported in seconds for Datasets 1 and 2.
6.3. Comparative Evaluation and Discussion
We evaluate our proposed clustering method against representative baseline methods, including EDR (Edit Distance on Real sequences), LCSS (Longest common subsequence), and DWT (Dynamic Time Warping), and the T2VEC deep learning model, selected from the benchmark study by Wang et al. (61) on Dataset 1. Additionally, we investigate the impact of the STC algorithm on all considered methods. Given that the dataset required preprocessing to fit our framework, we re-implemented the baselines to ensure a fair and consistent comparison.
Table 3 presents the runtime and performance results based on NMI and ARI metrics, where higher values indicate the efficiency of the STC algorithm in disregarding insignificant anomalies, directly resulting in better clustering performance. Based on the 19 ground-truth clusters established in the dataset, we optimized the DBSCAN hyperparameters ($\varepsilon$ and $MinLns$) to obtain 19 whole-trajectory clusters with the minimum possible number of outliers across all methods. The results show that the STC algorithm effectively reassigned outlier trajectories with only minor deviations to their appropriate clusters. Consequently, the NMI and ARI scores increased significantly across all methods.
Traditional methods rely on shape-based similarity between trajectories. Although DTW achieved the most accurate clustering among the baselines, it was computationally slower than our method. LCSS performed reasonably well but incurred the highest execution time, and EDR produced the weakest results. Thus, the similarity measure adopted in our work led to better clustering outcomes compared to the traditional approaches.
The deep learning model T2VEC achieved competitive results on the dataset. However, several limitations restrict its practicality. The model depends heavily on large training datasets (millions of trajectories) to extract meaningful trajectory representations. On small or medium-sized datasets, such as the one used in this study, the risk of underfitting and weak generalization increases. Another important limitation lies in interpretability. The model produces high-dimensional vector representations, but it is difficult to explain why two trajectories are considered similar. In contrast, the geometric principles used in our method are intuitive and directly reflect observable properties of the movement, such as curvature and overall shape. Furthermore, deep learning models are sensitive to a wide range of hyperparameters that are often complex to tune. Our approach depends on two parameters ($\varepsilon$ and $MinLns$) with clear geometric meaning, and provides consistent results across different runs. Finally, we could not apply the STC algorithm to T2VEC because of a fundamental difference in how they work. STC requires checking trajectory segments within specific time intervals. However, T2VEC converts the entire trajectory into a single vector, which hides the detailed time information necessary for this process. Although T2VEC performed competitively on the dataset, its strong results can be understood as an indirect approximation of the same geometric patterns that our method captures more simply and efficiently.
Category Method NMI ARI Time (s)
Traditional EDR 0.6521 0.3027 1526
EDR (STC) 0.7450 0.4618 8470
LCSS 0.7285 0.5043 8764
LCSS (STC) 0.8794 0.6311 16080
DTW 0.8622 0.7197 2524
DTW (STC) 0.9459 0.8359 12430
Deep Learning T2VEC 0.8098 0.7014 1380
T2VEC (STC) N/A N/A N/A
This Work Whole-Traj. 0.8542 0.6832 379
Whole-Traj. (STC) 0.9103 0.7340 912
Table 3. A performance comparison of clustering methods before and after the STC algorithm. The STC post-processing step shows a consistent boost in NMI and ARI scores for the traditional methods and our proposed approach. The WholeTraj.(STC)Whole-Traj.(STC) method indicates a strong balance between high accuracy and computational efficiency. Note: the STC is not applicable to T2VEC due to its non-geometric vectors; its execution time was measured on a GPU.
7. Conclusion
@@
Trajectory clustering algorithms can analyze either whole trajectories or subsets of them. In this paper, we propose both a whole-trajectory clustering algorithm and a sub-trajectory clustering algorithm. To discover the continuous object movements in each time interval, the split and merge procedures for line segments were implemented via DBSCAN line segment clustering. We used the average Euclidean distance to measure the distance between line segments. The result of line segment clustering is used not only for sliding-window sub-trajectory clustering to recognize local similarities, but also in whole-trajectory clustering to capture the similarity of object movements across all time intervals. Furthermore, we introduced a novel algorithm, called the stable trajectory clustering algorithm, that represents an advanced leap forward in the literature on trajectory clustering. This algorithm addresses scenarios where a moving object temporarily diverges from its cluster and then rejoins. Ignoring these minor anomalies at large scales helps analysts perform trajectory clustering without being distracted by insignificant deviations, maintains data integrity, and improves clustering accuracy. For example, vehicles stopping at traffic lights, pedestrians stopping at shop windows, or animals briefly detouring during migration should still be considered part of the main movement cluster. To determine whether the anomalies of outlier trajectories are real or temporary, the number of time intervals in which deviations occur, and the magnitudes of those deviations are important factors. Thus, we used the mean absolute deviation concept to analyze these factors. The empirical results on real data illustrated the effectiveness of the algorithm in detecting and ignoring insignificant anomalies. For this aim, NMI, ARI, and Silhouette score criteria were used to evaluate the quality of the clustering results before and after applying the stable trajectory clustering algorithm. The comparison showed that the stable clustering trajectory algorithm consistently increased the NMI and ARI scores. The results demonstrates its efficiency in handling insignificant anomalies to effectively improve clustering quality for both our proposed whole-trajectory method and traditional clustering methods. However, it was not applied to the deep learning method due to the non-geometric nature of its embeddings. The slight decrease in the Silhouette score is an expected consequence of this accuracy improvement. It confirms that the stable trajectory clustering algorithm focuses on creating semantically correct clusters, rather than ones that are merely geometrically compact.
The proposed algorithms exhibit sensitivity to the input parameters $\varepsilon$ and $MinLns$, with optimal values highly dependent on the characteristics of the moving data. This represents a limitation of our approach and warrants further investigation in future research. In addition, other ways to measure the stability of trajectory clustering can be used instead of the mean absolute deviation applied in this study. The definition of stability adopted in this work serves as the basis of our proposed algorithm. However, in other approaches, such as deep learning methods, it will be implemented differently.